OpenStudy (frostbite):

Need help solveing the inequality 4xy-3y^2>=0

OpenStudy (anonymous):

Hum, u have y(4x-3y) >=0 so seems you have cases...

OpenStudy (anonymous):

definitely cases. have two unknowns

OpenStudy (anonymous):

y>0 and y < 0 (y=0 as well...)

OpenStudy (anonymous):

are you solving for \(x\) in terms of \(y\)? because from what @estudier wrote, you see if \(y>0\) you need \(4x-3y>0\) making \(x>\frac{3}{4}y\) whereas if \(y<0\) it is the opposite

OpenStudy (frostbite):

Well i have just so far come to the conclution that x>= 3y/4 , y>=0 or x<=3y/4 ,y<=0

OpenStudy (frostbite):

Oh god why not i just write the full question i have been given: Determination of the domain and outline of the domain in an xy chart for the function f(x,y)=sqrt(4xy-3y^2)

OpenStudy (anonymous):

Yes, one is sort of reverse of the other....

OpenStudy (anonymous):

That should be a surface of some sort, hyperbolic maybe...

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=f%28x%2Cy%29+%3Dsqrt%284xy-3y^2%29 is what it looks like

OpenStudy (anonymous):

An infinite elliptic cone, lol

OpenStudy (frostbite):

But should it not be 2 dimensional when it is a xy-diagram?

OpenStudy (anonymous):

It is 2 dimensional, a surface...

OpenStudy (anonymous):

Not to mix up the dimensionality of the object and the space it is embedded in....

OpenStudy (anonymous):

U understand, right? f(x,y) is a function of 2 variables, so a surface...

OpenStudy (anonymous):

You may think of f(x,y) = z

OpenStudy (frostbite):

Yea beside i perhaps not made it clear that i am going to make a xy-diagram of the domain. i understand your point if you plot f(x,y), but that is not what i am going to find i think.